### v1

# Large Language Models Powered Context-aware Motion Prediction

Abstract
Motion prediction is among the most fundamental tasks in autonomous driving. Traditional methods of motion forecasting primarily encode vector information of maps and historical trajectory data of traffic participants, lacking a comprehensive understanding of overall traffic semantics, which in turn affects the performance of prediction tasks. In this paper, we utilized Large Language Models (LLMs) to enhance the global traffic context understanding for motion prediction tasks. We first conducted systematic prompt engineering, visualizing complex traffic environments and historical trajectory information of traffic participants into image prompts—Transportation Context Map (TC-Map), accompanied by corresponding text prompts. Through this approach, we obtained rich traffic context information from the LLM. By integrating this information into the motion prediction model, we demonstrate that such context can enhance the accuracy of motion predictions. Furthermore, considering the cost associated with LLMs, we propose a cost-effective deployment strategy: enhancing the accuracy of motion prediction tasks at scale with 0.7% LLM-augmented datasets. Our research offers valuable insights into enhancing the understanding of traffic scenes of LLMs and the motion prediction performance of autonomous driving.
I INTRODUCTION
Motion prediction is one of the most important tasks in the field of autonomous driving [1, 2, 3], which predicts the motion statuses of nearby agents by jointly considering nearby agents and road maps. This information will assist the decision module in making a more robust and safer driving decision. Therefore, this field already features a plethora of datasets and public competitions. For instance, since 2021, Waymo has been organizing competitions in motion prediction [^1] https://waymo.com/open/challenges/2021/motion-prediction/, attracting models and algorithms that have won championships, such as MTR [1], MTR++ [2], and MGTR [4]. There are three classes of motion prediction methods: goal-based methods [5, 6], direct-regression methods [7, 8], and methods that take the best of both[1, 2]. Although methods that take the best of both can balance results and computing resources greatly, they still encode information about agents and maps first and then decode to obtain final results. The methods above all lack a comprehensive understanding of overall traffic semantics.
Large Language Model (LLM) is popular in the field of autonomous driving, as not only the Language Model but also the World Model with common sense [9, 10, 11]. After the release of OpenAI’s GPT series of large language models, many works that use GPT to assist driving have been emerging [12, 13, 14, 15, 16]. These efforts leverage the inferential capabilities of large language models or through experimentation, aiding autonomous driving algorithms in better perception [17] and planning [12, 14, 15], and even facilitating end-to-end decision-making [18, 13]. However, teaching LLM to understand the interactive context among various transportation participants in diverse transportation scenarios is still a challenging task. Although many have applied vision-language models to traffic scenarios, their performance heavily relies on datasets annotated from a first-person perspective [19, 20, 21]. For abstract, information-overwhelming, and bird’s-eye view traffic scenarios, there is a lack of relevant datasets and understanding methodologies.
Figure 1: Context-aware Motion Prediction Based on LLMs. We first visualize structured information in the motion prediction dataset, GPT4-V then understand the scenario we provided via visualized image and well-designed prompt and output transportation context information. This information will be used to augment traditional motion prediction algorithms.
In this paper, we propose a new method that enhances the traffic context understanding of motion prediction models using LLMs to make more accurate predictions. To enable an unfine-tuned LLM—GPT4 with vision (GPT4-V) to understand traffic context and output the necessary context information for motion prediction, we visualize the vector map data and historical data of traffic participants as a Transportation Context Map (TC-Map) to serve as an image prompt and propose a corresponding text prompt design. After successfully obtaining traffic context information generated by the LLM, we integrated this information into a classical motion prediction algorithm—MTR [1]. The results indicate that this context information effectively improves the accuracy of motion prediction. Additionally, considering the cost of LLMs, we also propose a cost-effective deployment strategy: by utilizing 0.7% LLM-enhanced datasets, we can empower motion prediction performance at scale.
Our contributions are as follows:
  • We systematically designed and conducted prompt engineering to enable an unfine-tuned GPT4-V to comprehend complex traffic scenarios involving multiple participants and to output context information such as intention, affordance, and scenario;
  • We introduced a novel approach that combines the context information outputted by GPT4-V with the classical motion prediction pipeline [1], and we verified that this method enhances the effectiveness of motion prediction;
  • We proposed and validated a deployment strategy based on a dataset with a limited amount of LLM-generative context, which reduced the deployment cost of this method.
II Method
Figure 2: Details of Transportation Context Generation Prompt.
Figure 3: Six Prompt Design Guideline for GPT4-V Understanding Motion Predication Context.
II-A Get Transportation Context from GPT4-V
Translating images directly into motion poses is significantly challenging [22]. We obtain transportation context through methods based on LLMs with vision, defining the interested agent in the TC-Map as the ego-agent, which includes three categories: vehicle, pedestrian and cyclist. As shown in Fig. 2, the TC-Map designed based on the WOMD serves as the Rasterized Image prompt, and together with the Text prompt forms the Transportation Context Generation Prompt (TCGP). Through GPT4-V, this input yields outputs such as Situation Understanding, Reasoning, Intentions of ego-agent, Affordances [23], and Scenario Types. Additionally, during the prompt design process, we summarized 6 prompt design suggestions as shown in Fig. 3.
II-A1 TC-Map design suggestions
Corresponding to Prompt Design Suggestion 1 and 2 in Fig. 3. Suggestion 1 aims to reduce confusion in the volume of information. Inspired by the MTR Map Collection module, we introduced TC-Map Crop. Additionally, during the initial testing of GPT4-V, it was found that embedding a large amount of data information in the Text prompt, such as 11 frames of historical trajectory in JSON format, easily confuses GPT4-V and increases the uncertainty of intent. So the extent of TC-Map is cropped based on the type of agent (vehicle: pedestrian: cyclist = 6:4:3), limiting the ego-agent to focus only on the appropriate surrounding area scenes, reducing the impact of irrelevant factors on the prediction of Intention. Suggestion 2 aims to minimize confusion in direction identification. The diversity of scenarios leads to a variety of headings for the ego-agent, a common map guide icon is selected to normalize the heading of the ego-agent to “North”.
II-A2 Text Prompt design suggestions
Corresponding to Prompt Design Suggestion 3 and 4 in Fig. 3. Suggestion 3 proposes Examples, divided into Positive and Negative categories. Positive Examples adhere to the randomness and diversity of intent generation. Negative Examples are bad cases selected from GPT4-V’s responses during the quantitative experiment of prompts, focusing on scenarios with high misjudgment rates and uncommon, easily overlooked Corner-Cases. Additional Error Analysis is provided and placed closer to the end of the Text prompt. Suggestion 4 consists of two sentences. The first sentence discusses using rules to determine the lane type of the ego-agent’s driving lane in TC-Map, the second sentence provides explanations for the added lane types, they assist GPT4-V in excluding impossible intents under the premise that the ego-agent follows traffic rules, based on traffic driving regulations.
II-A3 Detailed Information Supplement
Corresponding to Prompt Design Suggestion 5 in Fig. 3. To fully leverage the scene information from WOMP dataset, such as road_edges and crosswalks, we optimize the representation of different types of scenario information in TC-Map by adding more indicative head arrows and visual labels for agents. During this phase, we also incorporate the speed of the ego-agent, as well as the speed, relative positions, and distances of the surrounding three agents.
II-A4 Multi-task Outputs Requirement
Corresponding to Prompt Design Suggestion 6 in Fig. 3, we designed seven types of Intentions for the ego-agent over the next 8 seconds. STRAIGHT-LEFT and STRAIGHT-RIGHT are considered scenarios where future ego-agent merging might occur, indicating complex intentions. Predicting these intentions often requires GPT4-V to utilize rasterized images, the ego-agent’s own position, heading, and speed, the state of surrounding vehicles, and the direction of nearby lane lines predicting these two types of intentions highly challenging. Additionally, based on an understanding of complex scenes, we designed five types of Affordances and four types of Scenario types, collectively providing a rich, human-like transportation context information to enhance the accuracy of the motion prediction task.
II-B Augment Motion Prediction Models via Transportation Context information
II-B1 Motion Transformer
We choose Motion TRansformer (MTR), SOTA of WOMD Challenge 2022, as our base model. MTR is developed based on the architecture of Transformer [24]. It first encodes the scenario information including agents and map roads, and then obtains the motion trajectories after several decoding processes.
II-B2 Integrate with Transportation Context Information
The transportation context contains three types of information: intention, affordance [23], and scenario in the format of a list, in which each element is a natural language description word. For example, [“STRAIGHT”, “RIGHT-TURN”], [“SLOW-ALLOW”, “RIGHT-ALLOW”], and [“INTERSECTION”] in Fig. 1, respectively. As the number of each type of context information is fixed, we use one-hot vectors to express the context information. In addition, because the description words in the list are not exclusive, we allow several ‘1’s to appear in the one-hot vector.
More specifically, for affordance and scenario information generated by LLM, we simply encode the nature language into one-hot vectors:
$A$ $=OneHotEncode(AffordanceList),$ (1)
$S$ $=OneHotEncode(ScenarioList),$ (2)
where $A\in\mathbb{R}^{8}$ and $S\in\mathbb{R}^{4}$ (see Fig. 1). As the number of ground truth intentions is only one, the number of description words in the intention information list is not strictly one. LLM gives possible intentions and the order of different words stands for different possibilities. So, when we encode intention information into a one-hot vector, we set different weights for different intentions in the list:
$I=WeightedOneHotEncode(IntentionList).$ (3)
The weight for i-th word in the list is:
$weight_{i}=length(IntentionInfoList)-i,$ (4)
where $I\in\mathbb{R}^{5}$ (see Fig. 1). So that the first word in the list has the maximal weight, and the last word holds the minimal weight.
To improve the performance of motion prediction and reduce the cost of computing resources, cluster-based intention points were provided in the process of decoding as prior knowledge [1]. Inspired by this, we also choose to integrate the transportation context information generated by LLM in the process of decoding.
The query content $Q\in\mathbb{R}^{K\times D}$ includes all scenario information in the data, where $K$ stands for the number of trajectories that need to be predicted and $D$ is the dimension of the feature. We first concatenate different types of context information and align the dimensions of traffic context information with query content, then integrate the context information generated by LLM via cross-attention after query content has been initialized.
$TC$ $=MLP([I,A,S]),$ (5)
$TC$ $=RepeatFirstDim(TC,K),$ (6)
$Q_{tc}$ $=CrossAttn(q=Q_{0},k=TC,v=TC),$ (7)
where $TC\in\mathbb{R}^{D}$ in Eq. 5 after multi-layer perceptron, and $TC\in\mathbb{R}^{K\times D}$ after repeating $K$ times in Eq. 6. When conducting cross-attention, we use initialized query content $Q_{0}$ as query, and $TC$ for both key and value in Eq. 7.
The following pipeline keeps all the same with MTR.
TABLE I: Effects of Different Components in Transportation Context Generation Prompt.
Sugg 1 Sugg 2 Sugg 3 Sugg 4 Sugg 5
TC-Map Crop DN NE+Corner-Case R-LT TCGP-Opt SPD-Add ACC(1st-$I$) ACC ACC(S,S-L,S-R)
0.3551 0.5140 0.5794
0.4579 0.5981 0.7383
0.4205 0.6635 0.7757
0.4112 0.6822 0.8037
0.5607 0.8317 0.9252
0.5514 0.8691 0.9252
0.5794 0.8691 0.9345
Sugg: Prompt Design Suggestions. TCGP-Opt: Coloring the agents surrounding the red ego-agent in different colors to distinguish them easily and optimizing the representation of scene elements in the TC-Map and supplementing the corresponding scene description in text prompts. SPD-Add: Adding information about the speed of the ego-agent, the speeds of surrounding agents, and their relative positions and distances to the ego-agent in text prompts. ACC: The accuracy of intentions output by GPT4-V. When GPT4-V outputs multiple intentions, if at least one intention matches the ground truth without additional clarification, it is considered correct. 1st-$I$: Taking only the first intention from the list of intentions output by GPT4-V. (S,S-L,S-R) means consolidating the three intentions of STRAIGHT, STRAIGHT-LEFT, and STRAIGHT-RIGHT into a single STRAIGHT category.
III Experiement
III-A Data distribution
Figure 4: Confusion Matrix of Intention Generated from GPT4-V
TABLE II: Data and Intention Accuracy of Three Ego-Agent Types.
method type ACC(1st-$I$) ACC ACC(S,S-L,S-R)
MTR V 0.8097 0.9442 0.9635
P 0.6906 0.9495 0.9564
C 0.7226 0.8822 0.9321
AVG 0.76448 0.93211 0.95326
GPT4-V V 0.6224 0.8482 0.9080
P 0.6222 0.8038 0.8133
C 0.5838 0.7857 0.9030
AVG 0.6145 0.8259 0.8865
We created a confusion matrix for the results in Tab. II as shown in Fig. 4, where we have merged STRAIGHT_LEFT and STRAIGHT_RIGHT into STRAIGHT. In this matrix, the rows represent the Ground Truth intentions, while the columns represent the intentions predicted by GPT4-V. Observing the main diagonal of the confusion matrix, it’s not hard to see that STRAIGHT has the highest prediction accuracy, followed by LEFT_TURN and RIGHT_TURN predictions. The lower prediction accuracy for STATIONARY is primarily because it can be correctly identified only when the ego-agent’s speed is zero. Nevertheless, its Ground Truth is classified as STATIONARY even at very low speeds. Ego-agents intended for LEFT_U_TURN are frequently classified as LEFT_TURN, a consequence of the WOMD map’s left-turn lanes occasionally accommodating straight movements, thus increasing the likelihood of LEFT_U_TURN and LEFT_TURN being predicted as STRAIGHT. Similarly, RIGHT_U_TURN ego-agents are often mistaken as STRAIGHT, and unexpectedly, they are frequently classified as STATIONARY. This anomaly primarily arises because intentions like RIGHT_U_TURN are most common among pedestrians, whose speed and walking direction can change abruptly, posing significant challenges to accurately predicting their intentions. We also present the results of MTR, even though the result of GPT-4V on Intention is lower than MTR, it can also provide information for motion prediction and improve the accuracy of MP as shown in Tab. III.
III-B Ablation Study of TCGP
We perform ablation studies on TCGB to understand the effectiveness of each component. As shown in Tab. I,
The data for the ablation experiment were randomly sourced from the WOMD testing set, comprising 107 TC-Maps samples. We have the following observations: 1) TCGP achieves optimal performance when fully equipped with all its components, with every module playing a role in predicting intentions. 2) TC-Map Crop enables GPT4-V to allocate more attention to the areas that should be focused on when predicting the future intentions of the ego-agent, significantly enhancing the accuracy of intention prediction. 3) R-LT is crucial because it provides the type of lane in which the ego-agent is located, and the type of lane often determines the future direction of the vehicle.
III-C LLM Augmented Motion Prediction
Dataset and metrics. We evaluate our methods and MTR on the large-scale WOMD [25]. For the task on WOMD, 1 second of history data will be provided, and 6 motion trajectories of the agents for 8 seconds in the future are needed. We utilize the official evaluation tool to calculate the evaluation metric, in which the mAP and miss rate (MR) are the most important ones in the official leaderboard  [^2] https://waymo.com/open/challenges/2023/motion-prediction/.
Data generation. There are roughly 2 million agents in WOMD and our prompt (text prompt and image prompt) costs about \$0.1 for each agent. So generating the transportation context information for all agents in WOMD costs nearly \$200,000. As it is costly to use GPT4-V for each agent in WOMD, we first generate the transportation context information from a minimal part of WOMD and then generate similar transportation context information for the remaining part via the nearest neighbor algorithm (see Alg.1). We first split the overall dataset into two different parts: $T_{1}$ and $T_{2}$, LLM will generate transportation context information for $T_{2}$. Secondly, we encode the scenario information for each agent in $T_{1}$ and $T_{2}$ and obtain their feature vectors. Finally, for each agent $A_{i}$ in $T_{1}$, the similar transportation context is the same as the agent $A_{j}$ in $T_{2}$ that has the closest distance with $A_{i}$. More specifically, we use Euclidean distance in the nearest neighbor finding algorithm.
Algorithm 1 Generate Scaled LLM Augmented Training Data with Minimal LLM Augmented Data
1:$T$: the overall dataset without LLM augmented; $Encoder$: scenario info encoder; $LLM$: large language model.
2:$T^{*}$: the overall dataset with LLM augmented.
3:initial $T_{1},T_{2}\leftarrow split(T)$, where $\left|T_{1}\right|\gg\left|T_{2}\right|$
4:$TC_{1}\leftarrow[]$, where $TC_{1}$ is Transportation Context info
5:$TC_{2}\leftarrow LLM(T_{2})$
6:$F_{1}\leftarrow Encoder(T_{1})$, where $F_{1}$ is feature vector list
7:$F_{2}\leftarrow Encoder(T_{2})$
8:for agent i in $T_{1}$ do
9:     $f_{i}\leftarrow F_{1}[i]$, where $f_{i}$ is a feature vector
10:     $j\leftarrow\textbf{NearestNeighbor}(f_{i},F_{2})$
11:     $TC_{1}[i]=TC_{2}[j]$
12:end for
13:$T^{*}_{1}\leftarrow Concatenate(T_{1},TC_{1})$
14:$T^{*}_{2}\leftarrow Concatenate(T_{2},TC_{2})$
15:$T^{*}\leftarrow Concatenate(T^{*}_{1},T^{*}_{2})$
Performance comparison. We compared our LLM augmented motion prediction model with MTR on the validation set of WOMD. We trained both models on 20% WOMD dataset. Our main results are in Table III. V, P, C are abbreviations of vehicle, pedestrian, cyclist, respectively. The result shows that our LLM augmented motion prediction model (with generated similar context information) outperformed MTR among all three agents (i.e. vehicle, pedestrian, and cyclist), and the average mAP increased by 0.95% compared with MTR.
TABLE III: Performance on the validation set of WOMD
[…]
MTR V 0.3862 0.8257 1.6789 0.1809
P 0.3587 0.3825 0.8088 0.0935
C 0.2848 0.7985 1.6380 0.2212
AVG 0.3432 0.6689 1.3752 0.1652
+LLM V 0.3954 0.8147 1.6205 0.1751
P 0.3754 0.3830 0.8070 0.0934
C 0.2924 0.8102 1.6464 0.2306
AVG 0.3527 0.6693 1.3580 0.1664
Performance Comparison. With LLM-augmented traffic context information, the motion prediction accuracy of different traffic agents improved. The best scores are expressed in bold.
Ablation Study. This study evaluates the contribution of each component within the transportation context generated by the LLM toward enhancing motion prediction accuracy. An ablation experiment was conducted by randomly selecting 5% of scenarios (approximately 24,000 scenarios) from the WOMD training set. The evaluation involved comparing the baseline model against the LLM-augmented model on the minimal part of WOMD validation set with traffic context information generated by LLM. The results are presented in Table IV. I, A, and S are abbreviations of intention, affordance, and scenario, respectively. V-mAP, P-mAP, and C-mAP mean mAP scores for vehicle, pedestrian, and cyclist, respectively. The results highlight the incremental value added by incorporating various types of transportation context information. Specifically, augmenting the model with all three types of transportation context information yielded the highest mAP score. Compared to the baseline MTR model, the final LLM-augmented model demonstrated a significant improvement, with a 1.67% increase in the mAP score.
TABLE IV: Ablation Experiment of
LLM Augmented Motion Prediction
method I A S mAP V-mAP P-mAP C-mAP
MTR - - - 0.2970 0.3300 0.3507 0.2102
- - - 0.3066 0.3377 0.3400 0.2422
- - - 0.2995 0.3386 0.3277 0.2322
- - - 0.3059 0.3341 0.3564 0.2272
- - 0.3030 0.3294 0.3458 0.2337
- - 0.3134 0.3318 0.3692 0.2392
- - 0.3117 0.3389 0.3554 0.2409
+LLM 0.3137 0.3385 0.3549 0.2476
Ablation Experiment. Results demonstrate that the model that uses three types of transportation context information gets the highest mAP score. The highest mAP score for each types of agent and average mAP score are expressed in bold.
IV Conclusions
In this paper, through systematic prompt engineering, we utilized the common sense and reasoning abilities of LLMs to extract human-like global context information from complex traffic scene motion predictions. By integrating these contexts into the traditional motion prediction pipeline, we enhanced the accuracy of trajectory prediction. We also proposed a cost-effective deployment strategy and verified its effectiveness.
Through meticulous prompt design, we achieved an impressive accuracy rate, underscoring the remarkable capacity of LLMs to comprehend complex and detailed transportation scenarios. To our knowledge, our work is pioneering in incorporating BEV-like TC-Maps into LLM prompts, offering a novel perspective for LLMs to interpret driving scenarios. Our experimental findings suggest that integrating transportation context information significantly enhances motion prediction capabilities. It is noteworthy that enabling LLMs to understand complex scene information remains challenging. Significant opportunities exist for exploring the integration of rich contextual information with motion prediction models. Our preliminary results suggest considerable potential for future development in this research area.

### v3

@@
-Large Language Models Powered Context-aware Motion Prediction
+Large Language Models Powered Context-aware Motion Prediction in Autonomous Driving
 Abstract
@@
 Our research offers valuable insights into enhancing the understanding of traffic scenes of LLMs and the motion prediction performance of autonomous driving.
+The source code is available at https://github.com/AIR-DISCOVER/LLM-Augmented-MTR and https://aistudio.baidu.com/projectdetail/7809548.
+Figure 1: Context-aware Motion Prediction Based on LLMs.
+We first visualize the structured information of one scenario in the motion prediction dataset, GPT4-V then understand the scenario via a visualized image and well-designed prompt.
+Finally, GPT4-V outputs transportation context information.
+This information will be used to augment traditional motion prediction algorithms.
 I INTRODUCTION
@@
 There are three classes of motion prediction methods: goal-based methods [5, 6], direct-regression methods [7, 8], and methods that take the best of both[1, 2].
-Although methods that take the best of both can balance results and computing resources greatly, they still encode information about agents and maps first and then decode to obtain final results.
+Although methods that take the best of both can balance results and computing resources greatly, they still encode information of agents and maps first and then decode to obtain final results.
 The methods above all lack a comprehensive understanding of overall traffic semantics.
@@
 Although many have applied vision-language models to traffic scenarios, their performance heavily relies on datasets annotated from a first-person perspective [19, 20, 21].
-For abstract, information-overwhelming, and bird’s-eye view traffic scenarios, there is a lack of relevant datasets and understanding methodologies.
-Figure 1: Context-aware Motion Prediction Based on LLMs.
-We first visualize structured information in the motion prediction dataset, GPT4-V then understand the scenario we provided via visualized image and well-designed prompt and output transportation context information.
-This information will be used to augment traditional motion prediction algorithms.
+For abstract, information-overwhelming, and bird’s-eye view (BEV) traffic scenarios, there is a lack of relevant datasets and understanding methodologies.
 In this paper, we propose a new method that enhances the traffic context understanding of motion prediction models using LLMs to make more accurate predictions.
-To enable an unfine-tuned LLM—GPT4 with vision (GPT4-V) to understand traffic context and output the necessary context information for motion prediction, we visualize the vector map data and historical data of traffic participants as a Transportation Context Map (TC-Map) to serve as an image prompt and propose a corresponding text prompt design.
+To enable an unfine-tuned LLM with vision (e.g.
+GPT4-V) to understand traffic context and output the necessary context information for motion prediction, we visualize the vector map data and historical data of traffic participants as a Transportation Context Map (TC-Map) to serve as an image prompt and propose a corresponding text prompt design.
 After successfully obtaining traffic context information generated by the LLM, we integrated this information into a classical motion prediction algorithm—MTR [1].
@@
 Our contributions are as follows:
-  • We systematically designed and conducted prompt engineering to enable an unfine-tuned GPT4-V to comprehend complex traffic scenarios involving multiple participants and to output context information such as intention, affordance, and scenario;
+  • We systematically designed and conducted prompt engineering to enable an unfine-tuned GPT4-V to comprehend complex traffic scenarios involving multiple traffic participants and to output context information such as intention, affordance, and scenario;
   • We introduced a novel approach that combines the context information outputted by GPT4-V with the classical motion prediction pipeline [1], and we verified that this method enhances the effectiveness of motion prediction;
   • We proposed and validated a deployment strategy based on a dataset with a limited amount of LLM-generative context, which reduced the deployment cost of this method.
+Figure 2: Details of Transportation Context Generation Prompt.
 II Method
-Figure 2: Details of Transportation Context Generation Prompt.
 Figure 3: Six Prompt Design Guideline for GPT4-V Understanding Motion Predication Context.
-II-A Get Transportation Context from GPT4-V
+II-A Get Transportation Context Information from GPT4-V
 Translating images directly into motion poses is significantly challenging [22].
-We obtain transportation context through methods based on LLMs with vision, defining the interested agent in the TC-Map as the ego-agent, which includes three categories: vehicle, pedestrian and cyclist.
-As shown in Fig.
-2, the TC-Map designed based on the WOMD serves as the Rasterized Image prompt, and together with the Text prompt forms the Transportation Context Generation Prompt (TCGP).
-Through GPT4-V, this input yields outputs such as Situation Understanding, Reasoning, Intentions of ego-agent, Affordances [23], and Scenario Types.
-Additionally, during the prompt design process, we summarized 6 prompt design suggestions as shown in Fig.
+We leverage vision-enhanced LLMs to derive transportation context, identifying the agent of interest in the TC-Map as the ego-agent, categorized into vehicles (V), pedestrians (P), and cyclists (C).
+Illustrated in Fig.
+2, the TC-Map designed based on the Waymo Open Motion Dataset (WOMD) [23] serves as the rasterized image prompt, and together with the text prompt forms the Transportation Context Generation Prompt (TCGP).
+The caption of the text prompt is obtained by filling in the template with code.
+Utilizing GPT4-V with this combined input facilitates the extraction of comprehensive context including situation understanding, reasoning, intentions of ego-agent, affordances [24], and scenario types.
+Additionally, six prompt design suggestions have been summarized for input prompts as shown in Fig.
 3.
 II-A1 TC-Map Design Suggestions
-Corresponding to Prompt Design Suggestion 1 and 2 in Fig.
-3.
-Suggestion 1 aims to reduce confusion in the volume of information.
-Inspired by the MTR Map Collection module, we introduced TC-Map Crop.
-Additionally, during the initial testing of GPT4-V, it was found that embedding a large amount of data information in the Text prompt, such as 11 frames of historical trajectory in JSON format, easily confuses GPT4-V and increases the uncertainty of intent.
-So the extent of TC-Map is cropped based on the type of agent (vehicle: pedestrian: cyclist = 6:4:3), limiting the ego-agent to focus only on the appropriate surrounding area scenes, reducing the impact of irrelevant factors on the prediction of Intention.
-Suggestion 2 aims to minimize confusion in direction identification.
-The diversity of scenarios leads to a variety of headings for the ego-agent, a common map guide icon is selected to normalize the heading of the ego-agent to “North”.
+Informed by prompt design suggestion 1 and 2.
+Suggestion 1 minimizes information overload.
+Drawing inspiration from the MTR Map Collection module [1], we introduced TC-Map Crop to focus the ego-agent’s attention solely on relevant surrounding scenes.
+Initial GPT4-V tests revealed that dense data, like 11 frames of historical trajectory in JSON format, leading to unclear intentions.
+Thus, we customized TC-Map dimensions: 120m x 120m for vehicles, 80m x 80m for pedestrians, and 60m x 60m for cyclists, respectively, to filter out extraneous influences on intention discernment.
+Furthermore, to address the varying orientations of the ego-agent across different scenarios, suggestion 2 proposes normalizing the heading of the ego-agent to “North” with a common map guide icon to reduce confusion in direction identification.
 II-A2 Text Prompt Design Suggestions
-Corresponding to Prompt Design Suggestion 3 and 4 in Fig.
-3.
-Suggestion 3 proposes Examples, divided into Positive and Negative categories.
-Positive Examples adhere to the randomness and diversity of intent generation.
-Negative Examples are bad cases selected from GPT4-V’s responses during the quantitative experiment of prompts, focusing on scenarios with high misjudgment rates and uncommon, easily overlooked Corner-Cases.
-Additional Error Analysis is provided and placed closer to the end of the Text prompt.
-Suggestion 4 consists of two sentences.
-The first sentence discusses using rules to determine the lane type of the ego-agent’s driving lane in TC-Map, the second sentence provides explanations for the added lane types, they assist GPT4-V in excluding impossible intents under the premise that the ego-agent follows traffic rules, based on traffic driving regulations.
+Addressing design suggestions 3 and 4, we introduce a structured approach to example creation.
+Suggestion 3 involves categorizing examples into positive and negative instances.
+Positive examples ensure a diverse and random generation of intentions.
+Negative examples are bad cases selected from GPT4-V’s responses during the quantitative experiment of prompts, focusing on scenarios with high misjudgment rates and uncommon, easily overlooked corner cases.
+Additional error analysis is provided and placed closer to the end of the negative example.
+Suggestion 4 consists of two sentences, the first sentence contains the ego-agent’s driving lane type in TC-Map determined by a rule-based algorithm, and the second sentence provides explanations for the added lane type.
+They aids GPT4-V in ruling out implausible intentions by assuming compliance with traffic regulations.
 II-A3 Detailed Information Supplement
-Corresponding to Prompt Design Suggestion 5 in Fig.
-3.
-To fully leverage the scene information from WOMP dataset, such as road_edges and crosswalks, we optimize the representation of different types of scenario information in TC-Map by adding more indicative head arrows and visual labels for agents.
-During this phase, we also incorporate the speed of the ego-agent, as well as the speed, relative positions, and distances of the surrounding three agents.
+Building on prompt design suggestion 5, we visualize more scenario elements using WOMD’s comprehensive scenario details, including road edges and crosswalks.
+For agents, we optimize the visualization of different scenario elements in the TC-Map by adding indicative head arrows and visual labels.
+Moreover, this suggestion includes integrating dynamic elements such as the ego-agent’s speed and the relative speeds, positions, and distances of nearby agents, enriching the context for more accurate motion prediction.
 II-A4 Multi-task Outputs Requirement
-Corresponding to Prompt Design Suggestion 6 in Fig.
-3, we designed seven types of Intentions for the ego-agent over the next 8 seconds.
-STRAIGHT-LEFT and STRAIGHT-RIGHT are considered scenarios where future ego-agent merging might occur, indicating complex intentions.
-Predicting these intentions often requires GPT4-V to utilize rasterized images, the ego-agent’s own position, heading, and speed, the state of surrounding vehicles, and the direction of nearby lane lines predicting these two types of intentions highly challenging.
-Additionally, based on an understanding of complex scenes, we designed five types of Affordances and four types of Scenario types, collectively providing a rich, human-like transportation context information to enhance the accuracy of the motion prediction task.
+Matching to prompt design suggestion 6, we delineated seven intention categories for the ego-agent inspired by the official evaluation tool of Waymo [23].
+STRAIGHT-LEFT and STRAIGHT-RIGHT anticipate potential lane changes, indicating the complexity of intentions.
+Predicting these intentions often requires GPT4-V to utilize rasterized images, the ego-agent’s position, heading, and speed, the state of surrounding vehicles, and the direction of nearby lane lines predicting these two intentions is highly challenging.
+Additionally, our approach includes the creation of five affordance types and four scenario types based on an understanding of comprehensive scenario analyses.
+These three types of information collectively provide rich, human-like transportation context information to enhance the accuracy of the motion prediction task.
 II-B Augment Motion Prediction Models via Transportation Context Information
 II-B1 Motion Transformer
-We choose Motion TRansformer (MTR), SOTA of WOMD Challenge 2022, as our base model.
-MTR is developed based on the architecture of Transformer [24].
-It first encodes the scenario information including agents and map roads, and then obtains the motion trajectories after several decoding processes.
+We selected the Motion Transformer (MTR) [1], the state-of-the-art model in the WOMD Challenge 2022, as our foundational model.
+Leveraging the Transformer architecture [25], MTR encodes comprehensive scenario data—including agent movements and map features—before decoding this information to predict motion trajectories.
+This process underpins our model’s ability to accurately anticipate future positions.
 II-B2 Integrate with Transportation Context Information
-The transportation context contains three types of information: intention, affordance [23], and scenario in the format of a list, in which each element is a natural language description word.
-For example, [“STRAIGHT”, “RIGHT-TURN”], [“SLOW-ALLOW”, “RIGHT-ALLOW”], and [“INTERSECTION”] in Fig.
-1, respectively.
-As the number of each type of context information is fixed, we use one-hot vectors to express the context information.
-In addition, because the description words in the list are not exclusive, we allow several ‘1’s to appear in the one-hot vector.
-More specifically, for affordance and scenario information generated by LLM, we simply encode the nature language into one-hot vectors:
+To enrich our motion prediction model with transportation context, we categorized context information into three types: intention, affordance, and scenario, each represented as natural language descriptors (e.g.
+[Straight, Right-Turn], [Slow-Allow, Right-Allow], and [Intersection] in Fig.
+1).
+These descriptors are encoded as one-hot vectors, allowing for multiple active elements to reflect the non-exclusive nature of these descriptors.
+Notably, affordances and scenarios are directly encoded, while intentions receive weighted encoding to prioritize them.
+More specifically, for affordance and scenario information generated by LLM, we simply encode the description words into one-hot vectors:
 $A$ $=OneHotEncode(AffordanceList),$ (1)
@@
 1).
-As the number of ground truth intentions is only one, the number of description words in the intention information list is not strictly one.
-LLM gives possible intentions and the order of different words stands for different possibilities.
+LLM gives all possible intentions and the order of different description words stands for different possibilities.
 So, when we encode intention information into a one-hot vector, we set different weights for different intentions in the list:
@@
 $weight_{i}=length(IntentionInfoList)-i,$ (4)
-where $I\in\mathbb{R}^{5}$ (see Fig. 1). So that the first word in the list has the maximal weight, and the last word holds the minimal weight.
+where $I\in\mathbb{R}^{5}$ (see Fig.
+1), and each intention’s weight inversely correlates with its position in the list, emphasizing more probable intentions.
+So that the first word in the list has the maximal weight, and the last word holds the minimal weight.
 To improve the performance of motion prediction and reduce the cost of computing resources, cluster-based intention points were provided in the process of decoding as prior knowledge [1].
@@
 The following pipeline keeps all the same with MTR.
-TABLE I: Effects of Different Components in Transportation Context Generation Prompt.
+TABLE I: Effects of Different Components in TCGP
 Sugg 1 Sugg 2 Sugg 3 Sugg 4 Sugg 5
@@
 0.5794 0.8691 0.9345
-Sugg: Prompt Design Suggestions.
-TCGP-Opt: Coloring the agents surrounding the red ego-agent in different colors to distinguish them easily and optimizing the representation of scene elements in the TC-Map and supplementing the corresponding scene description in text prompts.
-SPD-Add: Adding information about the speed of the ego-agent, the speeds of surrounding agents, and their relative positions and distances to the ego-agent in text prompts.
+Sugg: Prompt design suggestion.
+TCGP-Opt: Coloring agents surrounding the ego-agent in different colors to distinguish them easily, optimizing the visualization of scenario elements in the TC-Map and supplementing the corresponding scenario description in the text prompt.
+SPD-Add: Adding information about the speed of the ego-agent, the speeds of surrounding agents, and their relative positions and distances to the ego-agent in the text prompt.
 ACC: The accuracy of intentions output by GPT4-V.
 When GPT4-V outputs multiple intentions, if at least one intention matches the ground truth without additional clarification, it is considered correct.
-1st-$I$: Taking only the first intention from the list of intentions output by GPT4-V.
+(1st-$I$): Taking only the first intention from the list of intentions output by GPT4-V.
 (S,S-L,S-R) means consolidating the three intentions of STRAIGHT, STRAIGHT-LEFT, and STRAIGHT-RIGHT into a single STRAIGHT category.
 III Experiement
-III-A Data distribution
+III-A TCGP Generated Transportation Context Information
 Figure 4: Confusion Matrix of Intention Generated from GPT4-V
-TABLE II: Data and Intention Accuracy of Three Ego-Agent Types.
+TABLE II: Data and Intention Accuracy of Three Ego-Agent Types
 method type ACC(1st-$I$) ACC ACC(S,S-L,S-R)
@@
 C 0.7226 0.8822 0.9321
-AVG 0.76448 0.93211 0.95326
+AVG 0.7644 0.9321 0.9532
 GPT4-V V 0.6224 0.8482 0.9080
@@
 AVG 0.6145 0.8259 0.8865
-We created a confusion matrix for the results in Tab.
+Data distribution.
+The data set used to generate transportation context information was randomly selected from the WOMD validation set, consisting of approximately 15,000 TC-Map samples distributed according to a vehicle (V) to pedestrian (P) to cyclist (C) ratio of 9:3:3.
+We create a confusion matrix for the results of intention in Tab.
 II as shown in Fig.
-4, where we have merged STRAIGHT_LEFT and STRAIGHT_RIGHT into STRAIGHT.
-In this matrix, the rows represent the Ground Truth intentions, while the columns represent the intentions predicted by GPT4-V.
-Observing the main diagonal of the confusion matrix, it’s not hard to see that STRAIGHT has the highest prediction accuracy, followed by LEFT_TURN and RIGHT_TURN predictions.
-The lower prediction accuracy for STATIONARY is primarily because it can be correctly identified only when the ego-agent’s speed is zero.
-Nevertheless, its Ground Truth is classified as STATIONARY even at very low speeds.
-Ego-agents intended for LEFT_U_TURN are frequently classified as LEFT_TURN, a consequence of the WOMD map’s left-turn lanes occasionally accommodating straight movements, thus increasing the likelihood of LEFT_U_TURN and LEFT_TURN being predicted as STRAIGHT.
-Similarly, RIGHT_U_TURN ego-agents are often mistaken as STRAIGHT, and unexpectedly, they are frequently classified as STATIONARY.
-This anomaly primarily arises because intentions like RIGHT_U_TURN are most common among pedestrians, whose speed and walking direction can change abruptly, posing significant challenges to accurately predicting their intentions.
-We also present the results of MTR, even though the result of GPT-4V on Intention is lower than MTR, it can also provide information for motion prediction and improve the accuracy of MP as shown in Tab.
+4.
+For simplification, the categories STRAIGHT-LEFT and STRAIGHT-RIGHT were amalgamated into a single STRAIGHT category.
+The matrix’s rows denote the actual intentions, while its columns represent the intentions as predicted by GPT4-V.
+Observing the main diagonal of the confusion matrix, it’s not hard to see that STRAIGHT has the highest prediction accuracy, followed by LEFT-TURN and RIGHT-TURN predictions.
+The comparatively lower accuracy observed for STATIONARY intentions stems from its correct identification only at a zero speed of the ego-agent, despite being labeled as STATIONARY at minimal speeds.
+Ego-agents intended for LEFT-U-TURN are frequently classified as LEFT-TURN, a consequence of the WOMD map’s left-turn lanes occasionally accommodating straight movements, thus increasing the likelihood of LEFT-U-TURN and LEFT-TURN being predicted as STRAIGHT.
+Similarly, RIGHT-U-TURN ego-agents are often mistaken as STRAIGHT, and unexpectedly, they are frequently classified as STATIONARY.
+This anomaly primarily arises because intentions like RIGHT-U-TURN are most common among pedestrians, whose speed and walking direction can change abruptly, reflecting the challenge of precisely capturing the dynamic intentions of pedestrians.
+Despite GPT4-V’s intention prediction accuracy being lower than that of the MTR model, the integration of GPT4-V’s output still contributes valuable context for motion prediction, enhancing overall prediction accuracy, as evidenced in Tab.
 III.
-III-B Ablation Study of TCGP
-We perform ablation studies on TCGB to understand the effectiveness of each component. As shown in Tab. I,
-The data for the ablation experiment were randomly sourced from the WOMD testing set, comprising 107 TC-Maps samples.
+Ablation Study.
+We perform ablation studies on TCGP to understand the effectiveness of each component of our prompt in Tab.
+I.
+The data comprising 107 TC-Maps samples for the ablation experiment were randomly sourced from the validation set of WOMD.
 We have the following observations: 1) TCGP achieves optimal performance when fully equipped with all its components, with every module playing a role in predicting intentions.
 2) TC-Map Crop enables GPT4-V to allocate more attention to the areas that should be focused on when predicting the future intentions of the ego-agent, significantly enhancing the accuracy of intention prediction.
-3) R-LT is crucial because it provides the type of lane in which the ego-agent is located, and the type of lane often determines the future direction of the vehicle.
-III-C LLM Augmented Motion Prediction
+3) R-LT is crucial because it provides the type of lane in which the ego-agent is located, and the type of lane often determines the future direction of the ego-agent.
+These observations underscore the multifaceted contributions of TCGP’s components to the overall effectiveness of intention prediction, highlighting the importance of a holistic approach in optimizing predictive performance.
+III-B LLM Augmented Motion Prediction
 Dataset and metrics.
-We evaluate our methods and MTR on the large-scale WOMD [25].
-For the task on WOMD, 1 second of history data will be provided, and 6 motion trajectories of the agents for 8 seconds in the future are needed.
-We utilize the official evaluation tool to calculate the evaluation metric, in which the mAP and miss rate (MR) are the most important ones in the official leaderboard  [^2] https://waymo.com/open/challenges/2023/motion-prediction/.
+We evaluated our LLM-augmented motion prediction methods using the extensive WOMD dataset.
+Task requirements include predicting 6 future motion trajectories over 8 seconds, based on 1 second of historical data.
+The official WOMD evaluation tool calculates key metrics, notably the mean Average Precision (mAP), which serve as crucial benchmarks in the official leaderboard .
 Data generation.
-There are roughly 2 million agents in WOMD and our prompt (text prompt and image prompt) costs about \$0.1 for each agent.
-So generating the transportation context information for all agents in WOMD costs nearly \$200,000.
-As it is costly to use GPT4-V for each agent in WOMD, we first generate the transportation context information from a minimal part of WOMD and then generate similar transportation context information for the remaining part via the nearest neighbor algorithm (see Alg.1).
-We first split the overall dataset into two different parts: $T_{1}$ and $T_{2}$, LLM will generate transportation context information for $T_{2}$.
-Secondly, we encode the scenario information for each agent in $T_{1}$ and $T_{2}$ and obtain their feature vectors.
-Finally, for each agent $A_{i}$ in $T_{1}$, the similar transportation context is the same as the agent $A_{j}$ in $T_{2}$ that has the closest distance with $A_{i}$.
-More specifically, we use Euclidean distance in the nearest neighbor finding algorithm.
+To manage the high cost of generating transportation context information for WOMD’s 2 million agents (approximately \$0.1 per agent, totaling nearly \$200,000), we implemented a cost-effective strategy.
+Initially, LLM-generated transportation context for a small subset of WOMD was expanded to the remainder via a nearest neighbor algorithm, utilizing Euclidean distance for similarity assessment (see Alg.
+1).
+This approach efficiently scales LLM augmentation while maintaining data integrity.
 Algorithm 1 Generate Scaled LLM Augmented Training Data with Minimal LLM Augmented Data
@@
 We compared our LLM augmented motion prediction model with MTR on the validation set of WOMD.
-We trained both models on 20% WOMD dataset.
+We trained both models on the 20% WOMD dataset.
 Our main results are in Table III.
-V, P, C are abbreviations of vehicle, pedestrian, cyclist, respectively.
-The result shows that our LLM augmented motion prediction model (with generated similar context information) outperformed MTR among all three agents (i.e.
-vehicle, pedestrian, and cyclist), and the average mAP increased by 0.95% compared with MTR.
+Our LLM-augmented model exhibited superior performance over the MTR model across all agent types (vehicles, pedestrians, cyclists) on the WOMD validation set.
+Notably, the average mAP saw an enhancement of 0.95%, underscoring the augmented model’s effectiveness.
 TABLE III: Performance on the validation set of WOMD
@@
 AVG 0.3527 0.6693 1.3580 0.1664
-Performance Comparison. With LLM-augmented traffic context information, the motion prediction accuracy of different traffic agents improved. The best scores are expressed in bold.
+The best scores are expressed in bold.
 Ablation Study.
 This study evaluates the contribution of each component within the transportation context generated by the LLM toward enhancing motion prediction accuracy.
-An ablation experiment was conducted by randomly selecting 5% of scenarios (approximately 24,000 scenarios) from the WOMD training set.
-The evaluation involved comparing the baseline model against the LLM-augmented model on the minimal part of WOMD validation set with traffic context information generated by LLM.
+An ablation experiment was conducted by randomly selecting 5% of scenarios (approximately 24,000 scenarios) from the WOMD training set to train a model.
+The evaluation involved comparing the baseline model against the LLM-augmented model on the minimal part of the WOMD validation set with traffic context information generated by LLM.
 The results are presented in Table IV.
@@
 The results highlight the incremental value added by incorporating various types of transportation context information.
-Specifically, augmenting the model with all three types of transportation context information yielded the highest mAP score.
-Compared to the baseline MTR model, the final LLM-augmented model demonstrated a significant improvement, with a 1.67% increase in the mAP score.
+Incorporation of all context types resulted in the highest mAP scores, marking a 1.67% increase over the baseline MTR model, thereby validating the efficacy of LLM-augmented context in enhancing motion prediction.
 TABLE IV: Ablation Experiment of
 LLM Augmented Motion Prediction
-method I A S mAP V-mAP P-mAP C-mAP
+[…]
 MTR - - - 0.2970 0.3300 0.3507 0.2102
@@
 +LLM 0.3137 0.3385 0.3549 0.2476
-Ablation Experiment.
-Results demonstrate that the model that uses three types of transportation context information gets the highest mAP score.
-The highest mAP score for each types of agent and average mAP score are expressed in bold.
+The highest mAP score for each type of agent and the average mAP score is expressed in bold.
 IV Conclusions